# Representing Sugihara monoids

### v3

@@ 4.1 Closure of $\mathsf{RDInFL}$ under ultraproducts
Then $E_{\mathcal{F}}$ is an equivalence relation on $X$. Now consider the set of equivalence classes $Y=X/E_{\mathcal{F}}$. Define binary relations $\leqslant_{Y}$ and $E_{Y}$ on $Y$ by setting, for all $[x],[y]\in Y$,
@@
Suppose further that for each $i\in I$, the map $\alpha_{i}:X_{i}\to X_{i}$ is an order automorphism of $\mathbf{X}_{i}$ such that $\alpha_{i}\subseteq E_{i}$. Define $\alpha_{Y}:Y\to Y$ by setting, for each $[x]\in Y$,
@@
    If $0=\alpha_{Y}\circ\left(\leqslant_{Y}\right)^{c\smile}=\left(\leqslant_{Y}\right)^{c\smile}\circ\alpha_{Y}$, then $\left\langle\mathsf{Up}\left(\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle\right),\cap,\cup,\circ,\leqslant_{Y},0,\sim,-\right\rangle$ is a distributive InFL-algebra.
@@
We know that for each $i\in I$, there is a poset $\left\langle X_{i},\leqslant_{i}\right\rangle$, an equivalence relation $E_{i}$ on $X_{i}$ such that ${\leqslant_{i}}\subseteq E_{i}$, an order automorphism $\alpha_{i}$ of $\left\langle X_{i},\leqslant_{i}\right\rangle$ such that $\alpha_{i}\subseteq E_{i}$, and an embedding $\varphi_{i}$ from $\mathbf{A}_{i}$ into $\mathfrak{D}\left(\left\langle E_{i},\preccurlyeq_{i}\right\rangle\right)$. We may assume that $X_{i}\neq\varnothing$ for all $i\in I$ since the direct product of a family of algebras from $\mathsf{EDInFL}$ is isomorphic to another such product from which all factors of the form $\mathfrak{D}\left(\left\langle\varnothing,\varnothing\right\rangle\right)$ have been deleted. Now, for each $[a]\in A/\theta_{\mathcal{F}}$, define
@@
For any $[a]\in A/\theta_{\mathcal{F}}$, the set $R_{[a]}$ is an upset of $\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle$.
@@
Let $\left([x],[y]\right)\in R_{[a]}$ and assume $\left([x],[y]\right)\preccurlyeq_{Y}\left([u],[v]\right)$. The first statement implies that we have $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$. From the second part we get $[u]\leqslant_{Y}[x]$ and $[y]\leqslant_{Y}[v]$, which means $\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}$. Since $\mathcal{F}$ is closed under taking intersections, we have
@@
By Lemma 4.8 we know that $\varphi$ is well-defined. Now assume that $[a]\neq[b]$. This implies $(a,b)\notin\theta_{\mathcal{F}}$ and hence $\{i\in I\mid a(i)=b(i)\}\notin\mathcal{F}$. We get $\{i\in I\mid a(i)\neq b(i)\}\in\mathcal{F}$, and since each of the $\varphi_{i}$ is injective, it follows that $\{i\in I\mid a(i)\neq b(i)\}\subseteq\{i\in I\mid\varphi_{i}(a(i))\neq\varphi_{i}(b(i))\}$. But $\mathcal{F}$ is closed under taking supersets, so $\{i\in I\mid\varphi_{i}(a(i))\neq\varphi_{i}(b(i))\}\in\mathcal{F}$. Now
@@
We have to show that the ultraproduct $\mathbf{A}/\theta_{\mathcal{F}}$ of $\left\{\mathbf{A}_{i}\mid i\in I\right\}$ is representable. Recall that each of the $\mathbf{A}_{i}$ is embedded via $\varphi_{i}$ into an algebra of the form $\mathfrak{D}\left(\langle E_{i},\preccurlyeq_{i}\rangle\right)$. Each $\langle E_{i},\preccurlyeq_{i}\rangle$ comes from a poset $\mathbf{X}_{i}$ equipped with the necessary $E_{i}$ and $\alpha_{i}$. The definitions after Theorem 4.3 describe the structure $\langle E_{Y},\preccurlyeq_{Y}\rangle$, and, by Theorem 4.5(iv), $\left\langle\mathsf{Up}\left(\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle\right),\cap,\cup,\circ,\leqslant_{Y},\sim,-\right\rangle$ is a distributive InFL-algebra.
@@
To show that $\varphi$ preserves the monoid operation, first let $\left([x],[y]\right)\in\varphi\left([a]\right)\circ\varphi\left([b]\right)$. Then there is some $[z]\in Y$ such that $\left([x],[z]\right)\in\varphi\left([a]\right)$ and $\left([z],[y]\right)\in\varphi\left([b]\right)$. Hence, we obtain $\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$, which means
@@
The proof that $\varphi$ preserves -- is similar, and uses the definition of $\alpha_{Y}$.
@@
Let $\varphi:A\to B$ be defined by $\varphi(a)(i)=\varphi_{i}(a(i))$. Clearly $\varphi$ is a bijective map. Next, consider $\overline{\varphi}:A/\theta_{\mathcal{F}}\to B/\theta_{\mathcal{F}}$, defined by $\overline{\varphi}([a])=[\varphi(a)]$. We show that $\overline{\varphi}$ is bijective. For $[b]\in B/\theta_{\mathcal{F}}$, since $b\in B$, there exists $a\in A$ with $\varphi(a)=b$. So $\overline{\varphi}\left([a]\right)=\left[\varphi(a)\right]=[b]$. If $a,c\in A$ with $[a]\neq[c]$, then $(a,c)\notin\theta_{\mathcal{F}}$. Thus $\{\,i\in I\mid a(i)=c(i)\,\}\notin\mathcal{F}$, so $\{\,i\in I\mid a(i)\neq c(i)\,\}\in\mathcal{F}$. Since the $\varphi_{i}$ are injective, if $a(i)\neq c(i)$, then $\varphi_{i}(a(i))\neq\varphi_{i}(c(i))$ and so $\varphi(a)(i)\neq\varphi(c)(i)$. Hence $\{\,i\in I\mid a(i)\neq c(i)\,\}\subseteq\{\,i\in I\mid\varphi(a)(i)\neq\varphi(c)(i)\,\}$ so the latter set is in $\mathcal{F}$. Now $\{\,i\in I\mid\varphi(a)(i)=\varphi(c)(i)\,\}\notin\mathcal{F}$, so $(\varphi(a),\varphi(c))\notin\theta_{\mathcal{F}}$ and therefore $[\varphi(a)]\neq[\varphi(c)]$, i.e. $\overline{\varphi}\left([a]\right)\neq\overline{\varphi}\left([c]\right)$.
Next, we will show that $\overline{\varphi}$ preserves the fusion operation. For $[a],[c]\in A/\theta_{\mathcal{F}}$, we have
@@
Now consider the operation $\to^{\mathbf{A}/\theta_{\mathcal{F}}}$ on $\mathbf{A}/\theta_{\mathcal{F}}$. For $[a]$, $[c]\in A/\theta_{\mathcal{F}}$, we have
@@ 4.2 Representing all Sugihara monoids
The finitely generated subalgebras of $\mathbf{S}=(\mathbb{Z},\wedge,\vee,\cdot,\to,1,{\sim})$ are the algebras $\mathbf{S}_{2n+1}$ for $n\in\omega$. For the infinite Sugihara monoid $\mathbf{S}^{*}=(\mathbb{Z}{\setminus}\{0\},\wedge,\vee,\cdot,\to,1,{\sim})$, its finitely generated subalgebras are $\mathbf{S}_{2n+2}$ for $n\in\omega$. Lastly, recall that the quasivariety of Sugihara monoids is generated by $\{\mathbf{S},\mathbf{S}^{*}\}$ (cf. [17, Section 3]). We now have everything in place in order to prove our main result.
@@
[…]

### v4

@@ 4.1 Closure of $\mathsf{RDInFL}$ under ultraproducts
Then $E_{\mathcal{F}}$ is an equivalence relation on $X$. Now consider the set of equivalence classes $Y=X/E_{\mathcal{F}}$. Define binary relations $\leqslant_{Y}$ and $E_{Y}$ on $Y$ by setting, for all $x],[y]\in Y$,
@@
Suppose further that for each $i\in I$, the map $\alpha_{i}:X_{i}\to X_{i}$ is an order automorphism of $\mathbf{X}_{i}$ such that $\alpha_{i}\subseteq E_{i}$. Define $\alpha_{Y}:Y\to Y$ by setting, for each $x]\in Y$,
@@
[…]
@@
We know that for each $i\in I$, there is a poset $\left\langle X_{i},\leqslant_{i}\right\rangle$, an equivalence relation $E_{i}$ on $X_{i}$ such that ${\leqslant_{i}}\subseteq E_{i}$, an order automorphism $\alpha_{i}$ of $\left\langle X_{i},\leqslant_{i}\right\rangle$ such that $\alpha_{i}\subseteq E_{i}$, and an embedding $\varphi_{i}$ from $\mathbf{A}_{i}$ into $\mathfrak{D}\left(\left\langle E_{i},\preccurlyeq_{i}\right\rangle\right)$. We may assume that $X_{i}\neq\varnothing$ for all $i\in I$ since the direct product of a family of algebras from $\mathsf{EDInFL}$ is isomorphic to another such product from which all factors of the form $\mathfrak{D}\left(\left\langle\varnothing,\varnothing\right\rangle\right)$ have been deleted. Now, for each $a]\in A/\theta_{\mathcal{F}}$, define
@@
For any $a]\in A/\theta_{\mathcal{F}}$, the set $R_{[a]}$ is an upset of $\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle$.
@@
Let $\left([x],[y]\right)\in R_{[a]}$ and assume $\left([x],[y]\right)\preccurlyeq_{Y}\left([u],[v]\right)$. The first statement implies that we have $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$. From the second part we get $u]\leqslant_{Y}[x]$ and $y]\leqslant_{Y}[v]$, which means $\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}$. Since $\mathcal{F}$ is closed under taking intersections, we have
@@
By Lemma 4.8 we know that $\varphi$ is well-defined. Now assume that $a]\neq[b]$. This implies $(a,b)\notin\theta_{\mathcal{F}}$ and hence $\{i\in I\mid a(i)=b(i)\}\notin\mathcal{F}$. We get $\{i\in I\mid a(i)\neq b(i)\}\in\mathcal{F}$, and since each of the $\varphi_{i}$ is injective, it follows that $\{i\in I\mid a(i)\neq b(i)\}\subseteq\{i\in I\mid\varphi_{i}(a(i))\neq\varphi_{i}(b(i))\}$. But $\mathcal{F}$ is closed under taking supersets, so $\{i\in I\mid\varphi_{i}(a(i))\neq\varphi_{i}(b(i))\}\in\mathcal{F}$. Now
@@
[…]
@@
To show that $\varphi$ preserves the monoid operation, first let $\left([x],[y]\right)\in\varphi\left([a]\right)\circ\varphi\left([b]\right)$. Then there is some $z]\in Y$ such that $\left([x],[z]\right)\in\varphi\left([a]\right)$ and $\left([z],[y]\right)\in\varphi\left([b]\right)$. Hence, we obtain $\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$, which means
@@
The proof that $\varphi$ preserves - is similar, and uses the definition of $\alpha_{Y}$.
@@
Let $\varphi:A\to B$ be defined by $\varphi(a)(i)=\varphi_{i}(a(i))$. Clearly $\varphi$ is a bijective map. Next, consider $\overline{\varphi}:A/\theta_{\mathcal{F}}\to B/\theta_{\mathcal{F}}$, defined by $\overline{\varphi}([a])=[\varphi(a)]$. We show that $\overline{\varphi}$ is bijective. For $b]\in B/\theta_{\mathcal{F}}$, since $b\in B$, there exists $a\in A$ with $\varphi(a)=b$. So $\overline{\varphi}\left([a]\right)=\left[\varphi(a)\right]=[b]$. If $a,c\in A$ with $a]\neq[c]$, then $(a,c)\notin\theta_{\mathcal{F}}$. Thus $\{\,i\in I\mid a(i)=c(i)\,\}\notin\mathcal{F}$, so $\{\,i\in I\mid a(i)\neq c(i)\,\}\in\mathcal{F}$. Since the $\varphi_{i}$ are injective, if $a(i)\neq c(i)$, then $\varphi_{i}(a(i))\neq\varphi_{i}(c(i))$ and so $\varphi(a)(i)\neq\varphi(c)(i)$. Hence $\{\,i\in I\mid a(i)\neq c(i)\,\}\subseteq\{\,i\in I\mid\varphi(a)(i)\neq\varphi(c)(i)\,\}$ so the latter set is in $\mathcal{F}$. Now $\{\,i\in I\mid\varphi(a)(i)=\varphi(c)(i)\,\}\notin\mathcal{F}$, so $(\varphi(a),\varphi(c))\notin\theta_{\mathcal{F}}$ and therefore $\varphi(a)]\neq[\varphi(c)]$, i.e. $\overline{\varphi}\left([a]\right)\neq\overline{\varphi}\left([c]\right)$.
Next, we will show that $\overline{\varphi}$ preserves the fusion operation. For $a],[c]\in A/\theta_{\mathcal{F}}$, we have
@@
Now consider the operation $\to^{\mathbf{A}/\theta_{\mathcal{F}}}$ on $\mathbf{A}/\theta_{\mathcal{F}}$. For $[a]$, $c]\in A/\theta_{\mathcal{F}}$, we have
@@ 4.2 Representing all Sugihara monoids
The finitely generated subalgebras of $\mathbf{S}=\langle\mathbb{Z},\wedge,\vee,\cdot,\to,1,{\sim}\rangle$ are the algebras $\mathbf{S}_{2n+1}$ for $n\in\omega$. For the infinite Sugihara monoid $\mathbf{S}^{*}=\langle\mathbb{Z}{\setminus}\{0\},\wedge,\vee,\cdot,\to,1,{\sim}\rangle$, its finitely generated subalgebras are $\mathbf{S}_{2n+2}$ for $n\in\omega$. Lastly, recall that the quasivariety of Sugihara monoids is generated by $\{\mathbf{S},\mathbf{S}^{*}\}$ (cf. [17, Section 3]). We now have everything in place in order to prove our main result.
@@
Let $I=\{\,2k+1\mid k\in\omega\,\}$ and $\mathbf{A}=\prod\{\,\mathbf{S}_{i}\mid i\in I\,\}$. By Theorem 4.16, there exists an ultrafilter $\mathcal{F}$ on $I$ and an embedding $e_{1}:\mathbf{S}\hookrightarrow\mathbf{A}/\theta_{\mathcal{F}}$. Since $\mathbf{A}/\theta_{\mathcal{F}}\in\mathbb{P_{U}}(\mathsf{RSM})$ we get $\mathbf{S}\in\mathbb{ISP_{U}}(\mathsf{RSM})\subseteq\mathsf{RSM}$.
Next, let $J=\{\,2n+2\mid n\in\omega\,\}$ and $\mathbf{B}=\prod\{\,\mathbf{S}_{j}\mid j\in J\,\}$. By Theorem 4.16, there is an ultrafilter $\mathcal{G}$ on $J$ and an embedding $e_{2}:\mathbf{S}^{*}\hookrightarrow\mathbf{B}/\theta_{\mathcal{G}}$. Again, since $\mathbf{B}/\theta_{\mathcal{G}}\in\mathbb{P_{U}}(\mathsf{RSM})$ we get $\mathbf{S}^{*}\in\mathbb{ISP_{U}}(\mathsf{RSM})\subseteq\mathsf{RSM}$.